One number is 4 more than 2 times another. Their product is 30 . Find the numbers. Answer in the form of paired points with the lowest of the two numbers first.

Answer by TimothyLamb(4379)

---
x = 2y + 4
xy = 30
---
xy = 30
(2y + 4)y = 30
2yy + 4y - 30 = 0
---
the above quadratic equation is in standard form, with a=2, b=4, and c=-30
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 4 -30
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
y = 3
y = -5
---
this problem has two answers, the numbers are:
3, 10
or
-6, -5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php